OpenStudy (dls):

f(x)=log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10log10..............................n times.. What will be the domain of this function? (:

OpenStudy (anonymous):

Hmmm, interesting question...

OpenStudy (anonymous):

where is \(x\) though?

OpenStudy (unklerhaukus):

1

OpenStudy (anonymous):

1...what kind of question is this?...

OpenStudy (experimentx):

oh ,,, sorry, 1 produces complex values. try this \[ \huge [10^{10^{10^{10 ... \text{n-times}}}}, \infty)\]

OpenStudy (anonymous):

the domain would be \[-\infty \le x \le \infty\]

OpenStudy (experimentx):

only if \( f:\Bbb R\to \Bbb C \)

OpenStudy (anonymous):

f(x)=1...

OpenStudy (experimentx):

for that case you need to have the value 10^10^10^ ... n-1 times.

OpenStudy (experimentx):

woops!! EDIT:: \[ \huge \huge [10^{10^{10^{10 ... \text{(n-1) times}}}}, \infty)\]

OpenStudy (dls):

so what is the answer?